OpenStudy (anonymous):

Solve. 2x^2 + 9x = 35 x =5/2 and x = –7 x = 35 and x = 13 x = –5/2 and x = 7 x = –35 and x = –13

OpenStudy (anonymous):

Rewrite the problem as\[2x^{2}+9x-35=0\]then use the quadratic formula\[x=\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

Using a=2, b=9, c=-35

myininaya (myininaya):

\[2x^2+9x-35=0\] \[2x^2+14x-5x-35=0\] \[2x(x+7)-5(x+7)=0\] \[(x+7)(2x-5)=0\]

myininaya (myininaya):

x=-7, x=5/2

OpenStudy (anonymous):

FIne...if you want to factor it :)

OpenStudy (anonymous):

I swear I just did this one..

OpenStudy (anonymous):

Probably...people post and if they feel they don't get a response quickly, they repost the same problem.

OpenStudy (anonymous):

Haha posted a similar one :) sorry